Directions: In this activity you will be able to calculate the electric charges using the superposition principle. Solve the following problems. Show your solutions.



1. Three point charges, which are equally separated by a distance of 20cm, are arranged along the y-axis in a vacuum. The topmost charge bears a charge of -4.0μC, the middle charge has a charge of 3.0μC, and the bottom carries a 7.0μC. What is the magnitude and direction of the net electrostatic force that the middle charge experiences.



2. Find the force of attraction between two ball bearings with charges of 6μC and 14μC mm apart from each other.

1
Expert's answer
2022-01-25T06:51:34-0500

1. The magnitude of the force can be found by the superposition principle:


"F=\dfrac{q_2}{4\pi\epsilon_0r^2}(q_1+q_3)=2.02\text{ N}."


The force is directed upward.

2. The force of attraction between two ball bearings depends on the distance. It is unclear what the distance between them is, so, here is the equation, just plug in the distance in meters:


"F=\dfrac 1{4\pi\epsilon_0}\dfrac{q_1q_2}{r^2}=\dfrac{0.755}{r^2}\text{ N}."